OpenStudy (anonymous):

(verification) still typing .... Integral x+x^2tan(x) dx It is a definite integral from "-pi/2" to "pi/2", but I am just checking if I can find the general antiderivative of this function.

ganeshie8 (ganeshie8):

it looks odd

OpenStudy (anonymous):

\[\int\limits_{ }^{ } x+x^2\tan(x)~dx\]\[\frac{1}{2}x^2 + \int\limits_{ }^{ } x^2\tan(x)~dx\]\[\frac{1}{2}x^2 - x^2 \ln(\cos x) + \int\limits_{ }^{ } 2x \ln(\cos x)~dx\]so till now I am posting it.

OpenStudy (freckles):

when @ganeshie8 said it looks odd he meant not weird :p

OpenStudy (anonymous):

\[\int\limits_{ }^{ } \ln(\cos x ) dx\]\[x \ln(\cos x ) dx + \int\limits_{ }^{ }x \sin(x)\sec(x) dx\]\[x \ln(\cos x ) dx + \int\limits_{ }^{ }x \tan(x) dx\]\[x \ln(\cos x ) - x \ln(\cos x ) +\int\limits_{ }^{ } \ln(\cos x) dx\]so we have, \[\int\limits_{ }^{ } \ln(\cos x ) dx =x \ln(\cos x) -x \ln(\cos x) + \int\limits_{ }^{ } \ln(\cos x ) dx \]

OpenStudy (anonymous):

so the integral of ln(cosx) is zero? but that can't be.

OpenStudy (freckles):

you are doing a lot of work for nothing :p

OpenStudy (anonymous):

I just want to find the antiderivative, I know the rule for an odd function.

OpenStudy (freckles):

\[\text{ if }f(-x)=-f(x) \text{ then } F(-x)=F(x) \text{ where } F'=f \\ \int\limits_{-a}^{a}f(x) dx=F(x)|_{-a}^{a}=F(a)-F(-a)=F(a)-F(a)=0\]

OpenStudy (anonymous):

I probably did a mistake ln(cosx) dx in my second reply.

OpenStudy (anonymous):

reply 2, line 2 and 3, it is a minus, not a plus.

OpenStudy (anonymous):

I know knwo the odd function property. but I wanna solve it.

OpenStudy (anonymous):

So I'll start again from\[\frac{1}{2}x^2 - x^2\ln(\cos x) + \int\limits_{ }^{ } 2x \ln (\cos x ) dx \]

OpenStudy (freckles):

http://www.wolframalpha.com/input/?i=integrate%28x%5E2*tan%28x%29%2Cx%29 it doesn't look like an elementary integral

OpenStudy (anonymous):

actually, sorry, it is \[\int\limits_{ }^{ }x+x^2\tan(x)~dx~=\frac{1}{2}x^2 - x^2\ln(\cos x) + \int\limits_{ }^{ } 2x \ln (\cos x ) dx\]

OpenStudy (anonymous):

not an elemetary one, but I am just challenging myself. But perhaps you are right, and I'll just give up.

OpenStudy (anonymous):

Okay, I'll make it tan(x^3) lol.

OpenStudy (anonymous):

\[\int\limits_{ }^{ }x+x^2\tan(x^3) dx\]\[\frac{1}{2}x+\frac{1}{3}\int\limits_{ }^{ }\tan(u)~du\]\[\frac{1}{2}x+\frac{1}{3}\ln \left| x^3 \right|+C\]\[\frac{1}{2}x+\ln \left| x \right|+C\]

OpenStudy (anonymous):

but this is not fair though.

OpenStudy (anonymous):

nvm